OpenStudy (anonymous):

What is the solution to the inequality shown below? 7 3x is greater than -x + 5 A. x is less than 1 B. x is greater than 1 C. x is less than -1 D. x is greater than -1

OpenStudy (anonymous):

*** 7- 3 is greater than -x + 5

OpenStudy (anonymous):

Okay, which one? 7-3 > -x+5 7-3x > -x+5

OpenStudy (anonymous):

Craaaap. I got that wrong too. It's the second one.

OpenStudy (anonymous):

:-) Okay, so it's 7-3x > -x+5 First thing you want to do is set it up as an equation: 7-3x = -x+5 Then you want to get all the x terms on one side and the numbers on the other. Start by adding x to both sides: 7-3x+x = -x+x+5 = what?